# Training Language Models on Synthetic Edit Sequences Improves Code Synthesis

### v2

@@ 3.1 Pretraining Tiny LMs for Code Understanding
We begin our investigations by pre-training two tiny decoder-only transformers, TinyCodeLM-150M and TinyCodeLM-400M, for Python code understanding on 72 billion tokens of text. Pretraining our own language models grants us a data contamination-free test-bed to study code synthesis with edit sequences, rapidly evaluate LintSeq, and broadly re-examine the trade-off between inference-time compute and generation quality in code synthesis for models that can be updated on-device.
We rely on open-source data and libraries to pretrain our models (Penedo et al., 2024; Lozhkov et al., 2024; Soldaini et al., 2024; Groeneveld et al., 2024). Our pretraining data mix is inspired by Code Llama (Roziere et al., 2023), and reflects a code-skewed mixture of web text and raw Python sampled from FineWeb and TheStack, respectively (Penedo et al., 2024; Li et al., 2023). The architecture of our models respectively mimics the two smallest versions of GPT-2 (Radford et al., 2019), but integrates the transformer architecture changes proposed by the OLMo framework. This includes the absence of bias terms and the addition of non-parametric layer norms (Ba, 2016), as well as the use of SwiGLU (Shazeer, 2020), rotary positional embeddings (Su et al., 2024), and the GPT-NeoX-20B tokenizer (Black et al., 2022). We train both models for two epochs with a batch size of 524,288 tokens on an NVIDIA H100 node with four GPUs. Our experiments are supported by Pytorch FSDP (Zhao et al., 2023). More details on our pretraining procedures are in Appendix C.
@@ 3.2 Generating a Synthetic Dataset with LintSeq
Next, to support our finetuning experiments, we prepare a large “baseline” dataset of paired instruction and program data. We re-factorize the programs in this dataset into code edit sequences with LintSeq.
Table 1: Summary of temperature-tuned coding benchmark results for LMs with $\mathbf{\leq 0.2}$B parameters. Scores annotated with “($\dagger$)” indicate external model evaluations that we ran using the procedure described in Appendix B, and all other scores are as reported by model authors. We list models in order of increasing HumanEval pass@1 score.
@@
SmolLM-Instruct 135M 7.6 14.4($\dagger$) 10.1($\dagger$) 14.6($\dagger$) 🌑
TinyCodeLM-Instruct 150M 9.1 13.2 11.5 16.4 🌑
TinyCodeLM-LintSeqInstruct 150M 12.8 20.6 13.6 24.4 🌑
Table 2: Summary of temperature-tuned coding benchmark results for $\mathbf{0.2}$B to $\mathbf{0.4}$B parameter language models. Annotations, model order, and evaluation procedure are the same as in Table 1.
HumanEval MBPP(+)
Model Size pass@1 pass@10 pass@1 pass@10 Open-Source
PolyCoder 400M 3.0 5.3 - - 🌑
TinyCodeLM-Instruct 400M 9.6 18.5 15.5 22.2 🌑
SmolLM-Instruct 360M 11.3 19.3($\dagger$) 19.4($\dagger$) 23.1($\dagger$) 🌑
@@
Codegen-Mono 350M 12.8 23.1 9.4($\dagger$) 15.2($\dagger$) 🌑
@@
TinyCodeLM-LintSeqInstruct 400M 13.4 20.9 19.4 29.9 🌑
To that end, we first pool the Python portions of two open-source instruction datasets for code synthesis: the GPT 3.5/4-based Magicoder instruction tuning dataset and the StarCoder2-15B-based self-alignment training dataset (Wei et al., 2024b; a). These datasets are generated with the OSS-Instruct approach by Wei et al. (2024b) and have undergone decontamination for the benchmarks that we evaluate on in this paper. We conduct de-duplication on the pooled data to check for repeated examples. Furthermore, we strip any chain-of-thought-like natural language explanations from completion data. The resultant dataset has over 88,900 instruction-Python program pairs.
With our baseline dataset prepared, we run LintSeq to generate $s=5$ synthetic edit trajectory samples for each instruction-program pair. As described in Section 2.4, we concatenate each synthetic edit trajectory into a single string by interleaving consecutive edits with a special reserved “edit” token. Inspired by Muennighoff et al. (2024), we do not restrict against edit sequence repetitions. We use the popular Python linter pylint to guide edit sampling during generation. Examples of generated edit sequences and experiments testing the effect of varying $s$ are in Appendix E.
3.3 Finetuning Language Models on LintSeq Edit Sequences
We now probe the impact of instruction finetuning a variety of autoregressive LMs to synthesize code with edit sequences, compared to standard generation of full programs. We use two code synthesis benchmarks to support our model evaluations: HumanEval (Chen et al., 2021) and Mostly Basic Programming Problems (MBPP) (Austin et al., 2021).
Using both the code edit re-factorized and baseline instruction datasets obtained in section 3.2, we run pairs of finetuning experiments with six different models. In each experiment pair, we finetune an LM on both datasets for an equal number of optimizer steps and with the same learning rate schedule, saving intermediate checkpoints throughout finetuning. We run full benchmark evaluations on HumanEval and MBPP on each saved checkpoint [^1] To process the generations of edit sequence LMs into executable programs, we simply resolve each of the predicted code edits one-by-one. This procedure is visualized in Figure 1 and described in Appendix A.2., performing no prompt tuning. Then, we compare the best inference-time scaling behavior, i.e. benchmark coverage (pass@k) as a function of samples (k), obtained by edit sequence finetuning vs standard finetuning for each model. A more detailed description of the computed metrics as well as a full specification of the evaluation and finetuning procedures is provided in Appendices B and D.
Figure 4: HumanEval and MBPP coverage with repeated sampling (“pass@k vs. k”) achieved by instruction finetuning Gemma 2, Phi-3, and Llama 3.1 language models on a dataset of LintSeq edit sequence re-factored vs standard Python code (temperature $=1$, top-p $=0.95$).
@@ 3.3.1 TinyCodeLM
We run our first two pairs of finetuning experiments on TinyCodeLM-150M and TinyCodeLM-400M. These models were not pretrained on code synthesis instruction data, nor were they pretrained on any “diff”-like edit data. Our experimental results are summarized in Tables 1 and 2, where we compare the temperature-tuned performance of our models to the reported benchmark coverage of existing code LMs at similar parameter scales. We also report the inference-time scaling of benchmark coverage as a function of samples for our finetuned models in Appendix Tables 8 and 9.
For both the 150M and 400M parameter versions of TinyCodeLM, we find that finetuning LMs to synthesize code with edits via LintSeq data dramatically improves benchmark performance compared to baseline finetuning. Indeed, the edit sequence variants of TinyCodeLM outperform all existing code language models of comparable scale that we are aware of, including AlphaCode (Li et al., 2022b), Codex (Chen et al., 2021), CodeT5+ (Wang et al., 2023b), and the recent SmolLM-Instruct (Ben Allal et al., 2024). Our smaller edit sequence-finetuned model is particularly strong for its size, roughly matching or out-performing models with twice as many parameters including the 300M parameter version of Codex and the 302M-parameter version of AlphaCode (Tables 1 and 2).
@@
The results above raise a natural question: do performance improvements from finetuning LMs to synthesize code with edit sequences hold for other model scales, architectures, and tokenizers? To test this, we conduct four additional pairs of instruction finetuning experiments on LMs from three model families, Gemma 2, Phi-3, and Llama 3.1, ranging in size from 2.6B to 14B. We employ pretrained-only model weights, if available. The results of these experiments are in Figure 4, where we plot zero-shot benchmark coverage as a function of samples for instruction finetuned models. Raw coverage scores are reported in Appendix Tables 10 and 11.
Most of our findings echo those of Section 3.3.1. Aggregating across sample counts, we find that finetuning models to synthesize code with edits improves overall zero-shot performance on HumanEval and MBPP compared to finetuning on the original data. This suggests that re-factoring code with edit sequences is an architecture- and tokenizer-independent mechanism for improving downstream LM outputs. Furthermore, as shown in Figure 4 and Tables 10 and 11, we find that the degree by which edit sequence LMs outperform baseline model variants increases with repeated sampling for all tested models, culminating in an average absolute gain in pass@50 of +20% ($\pm$ 3%) on HumanEval and +12% ($\pm$ 2%) on MBPP. This observation confirms the hypothesis posed in Section 2, showing that training LMs to synthesize code with edits using LintSeq data improves the relationship between cumulative inference-time compute and zero-shot performance.
At pass@1, however, our results are slightly more mixed than in Section 3.3.1. For Phi-3 models, we observe either no difference or a decrease in score between each pair of model-data variants. One explanation for this is bias: the Phi-3 models have been previously instruction finetuned and were likely to have been trained on standard code data, putting LM variants finetuned to generate edit sequence re-factorized code at a comparative disadvantage due to distributional shift.
Figure 5: HumanEval and MBPP “pass@k vs. k” achieved by finetuning TinyCodeLM models on linter-guided vs randomly sampled code edit sequences (temperature $=1$, top-p $=0.95$).
3.4 Ablating Linter-Guidance
The backward sampling phase of LintSeq uses a linter or other verifier to group interdependent lines of a program together during edit sequence generation, ensuring that all code edits resolve to programs that are free of static errors. We conclude our experiments by testing the importance of this design choice with TinyCodeLM models.
@@ 3.4 Ablating Linter-Guidance
To do this, we replace the backwards procedure described in Section 2.2 with exclusively random sampling; during each step of the algorithm, we first sample the number of lines to delete from the current program uniformly at random, before sampling a set of lines with the desired count. Using this linter-ablated version of the algorithm, we generate a new synthetic edit sequence dataset with the same size as the LintSeq dataset used in all previous finetuning experiments, i.e. with $s=5$ example sequences per sample in the source dataset. The average number of edits per example in this dataset ($\overline{E}_{\text{RandSeqInstruct}}=3.9$) is empirically similar to its linter-guided counterpart ($\overline{E}_{\text{LintSeqInstruct}}=3.8$, see Figure 3). We employ the same procedure as the one used in Section 3.3 to instruction finetune TinyCodeLM models on the dataset of randomly sampled edit sequences.
In Figure 5, we compare the temperature $1$ inference-time scaling laws on HumanEval and MBPP obtained by finetuning models on randomly sampled vs static error-free edit sequences. Raw model scores are also provided in Appendix F, Tables 8 and 9. Ablating linter-guidance results in a decline in benchmark coverage. On HumanEval, linter ablation reduces absolute pass@50 score on TinyCodeLM-150M ($22.6\%\mapsto 17.7\%$) and on TinyCodeLM-400M ($26.8\%\mapsto 22.0\%$). MBPP pass@50 is similarly affected, dropping for both models ($34.5\%\to 30.2\%$ and $39.6\%\mapsto 34.5\%$). These results suggest that the error-free nature of edits in LintSeq instruction finetuning data does indeed have a positive impact on the coding problem coverage of sampled solutions.
To conclude our analysis, we probe whether training models on error-free edit sequences also has an effect on the presence of errors in across all generated programs, aside from its positive effects on coverage. To assess this, we run the Python linter pylint over each of the synthesized programs used to compute the reported temperature $1$ pass@k metrics, checking code for static errors.
@@
In Figure 6, we plot the total proportions of synthesized program samples with at least one static error across finetuned model variants. On both benchmarks, LMs trained on randomly sampled edits (dark grey) appear to generate “buggy” code with much higher frequency than all other models. Furthermore, on HumanEval, we find that LintSeq models (indigo) synthesize programs with static errors at a higher frequency than baseline models (light grey), despite their higher coverage of benchmark coding problems. This additional finding suggests that model performance gains from LintSeq cannot simply be attributed to improvement in static error frequency across code – training on re-factored code must be helping models write generally better, more diverse programs.
In summary, the error-free nature of the linter-guided edits sampled in LintSeq appears to indeed be important for improving both the quality and diversity of sampled programs (Figure 5), as well as the overall correctness (Figure 6) of code synthesized by language models trained on edit sequences.
Figure 6: Comparing static error frequency in synthesized code samples across baseline vs edit sequence instruction finetuned model variants (n $=50$, temperature $=1$, top-p $=0.95$).

### v3

@@ 3.1 Pretraining Tiny LMs for Code Understanding
We begin our investigations by pre-training two tiny decoder-only transformers, TinyCodeLM-150M and TinyCodeLM-400M, for Python code understanding on 72 billion tokens of text. Pretraining our own language models grants us a data contamination-free test-bed to study code synthesis with edit sequences, rapidly evaluate LintSeq, and broadly re-examine the trade-off between test-time compute and generation quality in code synthesis for models that can be updated on-device.
We rely on open-source data and libraries to pretrain our models (Penedo et al., 2024; Lozhkov et al., 2024; Soldaini et al., 2024; Groeneveld et al., 2024). Our pretraining data mix is inspired by Code Llama (Roziere et al., 2023), and reflects a code-skewed mixture of web text and raw Python sampled from FineWeb and The Stack, respectively (Penedo et al., 2024; Li et al., 2023). The architecture of our models respectively mimics the two smallest versions of GPT-2 (Radford et al., 2019), but integrates the transformer architecture changes proposed by the OLMo framework. This includes the absence of bias terms and the addition of non-parametric layer norms (Ba, 2016), as well as the use of SwiGLU (Shazeer, 2020), rotary positional embeddings (Su et al., 2024), and the GPT-NeoX-20B tokenizer (Black et al., 2022). We train both models for two epochs with a batch size of 524,288 tokens on an NVIDIA H100 node with four GPUs. Our experiments are supported by Pytorch FSDP (Zhao et al., 2023). More details on our pretraining procedures are in Appendix D.
@@ 3.2 Generating a Synthetic Dataset with LintSeq
To support our fine-tuning experiments, we prepare a baseline dataset of paired instruction and program data. We then re-express the programs in this dataset as code edit sequences with LintSeq.
To that end, we first pool the Python portions of two open-source instruction datasets for code synthesis: the GPT 3.5/4-based Magicoder instruction dataset and the StarCoder2-15B-based self-alignment training dataset (Wei et al., 2024b; a). These datasets are generated with the OSS-Instruct approach by Wei et al. (2024b) and have undergone decontamination for the benchmarks that we evaluate on in this paper. We conduct de-duplication on the pooled data to check for repeated examples. Furthermore, we strip any chain-of-thought-like natural language explanations from completion data. The resultant dataset has over 88,900 instruction+program pairs.
Table 1: HumanEval and MBPP(+) results for TinyCodeLMs after SFT vs existing code models of similar scale ($\leq$ 0.4B parameters). Scores annotated with $\dagger$” indicate external model evaluations that we ran using the procedure described in Appendix C, and all other scores are as reported by model authors. We list models in order of increasing HumanEval pass@1 and report standard error in computed score. Sampling hyperparameters are listed in Appendix C.4.
@@
SmolLM-Instruct 135M 7.7 $\pm$ 0.8$\dagger$ 14.5 $\pm$ 1.0$\dagger$ 10.1 $\pm$ 1.8$\dagger$ 14.6 $\pm$ 0.5$\dagger$ 🌑
TinyCodeLM-Instruct 150M 9.1 $\pm$ 2.3 13.5 $\pm$ 0.6 11.5 $\pm$ 1.9 21.6 $\pm$ 0.4 🌑
TinyCodeLM-Instruct 400M 11.3 $\pm$ 0.9 18.5 $\pm$ 1.1 15.5 $\pm$ 2.1 22.2 $\pm$ 0.5 🌑
SmolLM-Instruct 360M 11.3 19.3 $\pm$ 1.1$\dagger$ 19.4 $\pm$ 2.4$\dagger$ 23.1 $\pm$ 0.5$\dagger$ 🌑
@@
TinyCodeLM-LintSeqInstruct 150M 12.8 $\pm$ 2.6 20.6 $\pm$ 1.1 13.6 $\pm$ 2.1 24.4 $\pm$ 0.8 🌑
Codegen-Mono 350M 12.8 23.1 9.4 $\pm$ 1.8$\dagger$ 15.2 $\pm$ 0.7$\dagger$ 🌑
@@
TinyCodeLM-LintSeqInstruct 400M 13.4 $\pm$ 2.0 20.9 $\pm$ 1.1 19.4 $\pm$ 2.4 29.9 $\pm$ 0.6 🌑
With our baseline dataset prepared, we run LintSeq to generate $s=5$ synthetic edit sequences for each instruction-program pair. As described in Section 2.5, we concatenate each synthetic edit sequence into a single string by interleaving consecutive edits with a special reserved “edit” token. Inspired by Muennighoff et al. (2024), we do not restrict against edit sequence repetitions. We use the popular Python linter pylint to guide edit sampling during generation. Examples of generated edit sequences and experiments testing the effect of varying $s$ are in Appendix F.
3.3 Training Language Models on LintSeq Edit Sequences with SFT
Next, we probe the impact of training autoregressive LMs to synthesize full programs vs. program edit sequences according to natural language instructions. Aside from the tiny code LMs described above in Section 3.3.1, we also finetune small LMs from three different model families, ranging in scale from 2.6B to 14B parameters. We evaluate tiny code LMs on HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), and small LMs on the additional challenging benchmarks DS-1000 (Lai et al., 2023), BigCodeBench (Zhuo et al., 2024), and CodeContests (Li et al., 2022b).
Using both the refactored and baseline instruction datasets described in section 3.2, we run pairs of SFT experiments with six different models. In each experiment pair, we finetune an LM on both datasets for an equal number of optimizer steps and with the same learning rate schedule, saving intermediate checkpoints throughout fine-tuning. Then, we compare the benchmark performance of checkpoints across sampling temperatures [^2] To process the generations of edit sequence LMs into executable programs, we simply resolve each of the predicted code edits one-by-one. This procedure is visualized in Figure 1 and described in Appendix B.2., performing no prompt tuning. A more detailed description of the computed metrics as well as a full specification of the evaluation and fine-tuning procedures is provided in Appendices C and E.
@@ 3.3.1 TinyCodeLM
We run our first two pairs of fine-tuning experiments on TinyCodeLM-150M and TinyCodeLM-400M. Our experimental results are summarized in Table 1, where we compare the temperature-tuned performance of our models on HumanEval and MBPP(+) to the pass@1 and pass@10 scores of existing LMs with similar parameter counts.
For both the 150M and 400M parameter versions of TinyCodeLM, we find that fine-tuning LMs to synthesize code with edits via LintSeq data results in stronger benchmark performance compared to the baseline, improving HumanEval pass@1 by 41% ($9.1\mapsto 12.8$) and 19% ($11.3\mapsto 13.4$) and MBPP pass@1 by 18% ($11.5\mapsto 13.6$) and 25% ($15.5\mapsto 19.4$). We see a similar scale of improvement on pass@10 for both benchmarks. Our smaller LintSeq model is particularly strong for its size, roughly matching the performance of several models with larger parameter counts (Table 1).
Figure 3: HumanEval, MBPP(+), DS-1000, and BigCodeBench (Instruct) results for Gemma 2, Phi-3, and Llama 3.1 models after SFT on LintSeq (indigo) vs standard Python code (grey). On HumanEval and MBPP(+), we tune sampling temp., top-p, and min-p over $\{1,1.1,1.2\}$, $\{0.95,1.0\}$, and $\{0,0.05\}$, respectively with $n=64$ samples. On DS-1000, we evaluate models with the completion format, temperature $=0.2$, top-p $=0.5$, min-p $=0$, and $n=40$, following Wei et al. (2024b) and Luo et al. (2023). On BigCodeBench Instruct, we evaluate with greedy decoding (Zhuo et al., 2024). Error bars on HumanEval and MBPP scores show standard error.
@@
The results above raise a few questions: Do performance improvements from fine-tuning LMs to synthesize code with edit sequences also hold for language models that were not specifically pretrained for code understanding? Do they hold across model scales, architectures, and tokenizers?
To answer these questions, we conduct four additional pairs of SFT experiments on LMs from three model families, Gemma 2, Phi-3, and Llama 3.1. We use pretrained-only model weights, if available. The selected LMs range in size from 2.6B to 14B and were trained on general-purpose data mixtures (Gemma Team et al., 2024; Abdin et al., 2024; Dubey et al., 2024).
Our findings align with those presented in Section 3.3.1. As shown in Figure 3, LintSeq improves performance on each LMs for all but two of the metrics visualized here (HumanEval pass@1 and BigCodeBench Instruct greedy pass@1). Notably, even on these metric, the least performant LintSeq instruction-tuned models still achieve performance that is comparable to the baseline, i.e. within standard error of sampling or within a percentage point. In aggregate across models, LintSeq improves HumanEval, MBPP, DS-1000, and BigCodeBench Instruct pass@1 by an average absolute gain of +2.3, +4.3, +3.1, and +1.1 in score compared to baseline SFT.
Furthermore, as shown in Figure 1(right) and Figure 4, the degree by which edit sequence LMs outperform baselines on HumanEval, MBPP, and CodeContests increases with repeated sampling for all tested models. In each of the plots included in these figures, we show the total proportion of benchmark problems solved by SFT-ed LMs on any attempt given “k” tries as a function of total test-time compute used during repeated sampling. By comparing total test-time compute across model variants, we account for the slight difference between LintSeqInstruct vs Instruct model generation lengths due to the extra “diff” descriptor tokens used by edit sequence models. Even after adjusting for these extra tokens, LintSeq consistently improves the relationship between total test-time compute and performance on code synthesis, supporting the hypothesis posed in Section 2.
In summary, the results of these experiments suggest that refactoring code tuning data into synthetic edit sequences with LintSeq is a code-pretraining-, scale-, architecture-, and tokenizer-independent mechanism for improving the quality and diversity of LM outputs on code generation tasks.
3.4 Ablating the Linter from LintSeq
The backward sampling phase of LintSeq uses a linter to decompose code across edits whose contents reflect the syntactical structure of its programming language. We conclude our experiments by testing the importance of this design choice with TinyCodeLM models: does fine-tuning on sequences of (entirely) randomly sampled code edits hurt model performance on HumanEval and MBPP(+)?
Figure 4: Repeatedly sampling from models SFT-ed to generate edit seqs. vs full programs: we compare the best pass@k score achieved by modulating sampling hyperparameters for LintSeqInstruct vs Instruct models. On HumanEval and MBPP(+), we use the same values as in Figure 3, while on CodeContests, we sweep over temperatures $\{0.5,0.6\}$ and use top-p $=1.0$, min-p $=0$, and $n=128$. We then plot benchmark score as a function of the total cost of repeated sampling from each model in FLOPs (see Appendix A.4). Shading shows standard error in linear fit. See Figure 1 for Phi-3 3.8B and Llama 3.1 8B test-time scaling with repeated sampling curves on HumanEval and MBPP.
@@ 3.4 Ablating the Linter from LintSeq
To test this, we replace the backwards procedure described in Section 2.3 with fully random sampling; during each step of the algorithm, we first sample the number of lines to delete from the current program uniformly at random, before sampling a set of lines with the desired count. We refer to this algorithm as “RandSeq.” Using RandSeq, we generate a new synthetic edit sequence dataset with the same size as the LintSeq dataset used in all previous fine-tuning experiments. The average number of edits per example in this dataset ($\approx 3.9$) is similar to its linter-guided counterpart ($\approx 3.8$) [^3] Note that both datasets also have a similar size in total training tokens ($\approx 18\cdot 10^{6}$ TinyCodeLM tokens)..
We employ the same procedure as the one used in Section 3.3 to SFT TinyCodeLM models on the RandSeq dataset. In Figure 5(left), we compare the pass@1 HumanEval and MBPP score of LintSeqInstruct vs RandSeqInstruct models at high temperatures. On both benchmarks and models, ablating the linter from LintSeq hurts performance with statistical significance, reducing HumanEval pass@1 by 30% ($6.4\mapsto 4.5$) and 29% ($8.4\mapsto 6.0$) and MBPP pass@1 by 24% ($8.6\mapsto 6.5$) and 28% ($14.2\mapsto 10.2$), respectively. These results suggest that the linter-informed structure of edits in LintSeq fine-tuning data does improve model performance.
@@
In Figure 5(right), we conclude our analysis by probing whether training models on linted edits has an effect on the total proportion of syntactical errors in completed programs. To assess this, we run the Python linter pylint over the full set of generations sampled at temperature $=1$, top-p $=1$, and min-p $=0$, checking each generated program for syntax errors with this linter. LMs trained on randomly sampled edits appear to generate “buggy” code with much higher frequency than all other models on both HumanEval and MBPP(+). Furthermore, on HumanEval, we find that LintSeq models synthesize programs with linter-errors at a higher frequency than baselines, despite their higher pass@1. This additional finding suggests that model performance gains from LintSeq cannot simply be attributed to improvement in low-level correctness of generated code – training on refactored code must be helping models write generally better, more diverse programs.
Figure 5: Left: HumanEval and MBPP(+) pass@1 achieved by fine-tuning TinyCodeLM models on linter-guided (LintSeq) vs randomly sampled (RandSeq) code edit sequences. We tune sampling parameters over the same values as in Figures 3 and 4, and report the best scores for each model. Right: Comparing total proportions of generations with lint errors. Error bars show standard error.